### v1

# Generalizable Trajectory Prediction via Inverse Reinforcement Learning with Mamba-Graph Architecture

Abstract
Accurate driving behavior modeling is fundamental to safe and efficient trajectory prediction, yet remains challenging in complex traffic scenarios. This paper presents a novel Inverse Reinforcement Learning (IRL) framework that captures human-like decision-making by inferring diverse reward functions, enabling robust cross-scenario adaptability. The learned reward function is utilized to maximize the likelihood of output by the encoder-decoder architecture that combines Mamba blocks for efficient long-sequence dependency modeling with graph attention networks to encode spatial interactions among traffic agents. Comprehensive evaluations on urban intersections and roundabouts demonstrate that the proposed method not only outperforms various popular approaches in prediction accuracy but also achieves 2 times higher generalization performance to unseen scenarios compared to other IRL-based method.
I Introduction
Modern intelligent transportation systems (ITS) rely on accurate trajectory prediction to enhance road safety and traffic efficiency [1]. However, predicting vehicle trajectories in complex urban environments remains a significant challenge due to the intricate interplay of factors, including dynamic road geometries, dense traffic interactions, and the inherent stochasticity of human driving behavior [2].
Traditional prediction models with supervise learning paradigms aim to learn the mapping from historical trajectories to future trajectories, as well as the hidden state of the environment. Sequence modeling approaches, such as recurrent neural networks (RNNs) [3] and long short-term memory (LSTM) networks [4, 5, 6], have been widely employed to model temporal dependencies in trajectory data. With the advent of Mamba [7], a selective State Space Model (SSM), it offers superior capability in modeling long-range dependencies while maintaining exceptional inference efficiency—a critical attribute for real-world autonomous driving applications where computational latency is non-negligible. However, these methods often fall short in accurately capturing the intricate spatiotemporal interactions between vehicles and their environments. Graph-based encoding methods, such as Graph Neural Networks (GNNs) [8], have emerged as a promising solution to address the interrelationships between vehicles and their surroundings [9, 10, 11]. Yet, these models often struggle to capture the nuanced, context-dependent decision-making processes of human drivers, leading to suboptimal or unsafe predictions, particularly in scenarios involving complex maneuvers or unexpected events. Thus in this paper, we combined Mamba block and GNN in our predictor module.
To further address the challenges of dataset dependency and cross-scenario generalization, Inverse Reinforcement Learning (IRL) [12] has shown promise in trajectory prediction complementing these supervised learning approaches. Whereas pure IRL approaches [13, 14], constrained by their reliance on environment dynamics modeling, are only viable in simplistic settings, our IRL-augmented data-driven approach not only extends applicability to complex environments, but also preserves high predictive accuracy. Diverging from conventional paradigms [15, 16]—which hinge on maximum margin theory and iteratively refine policies through learner-expert performance comparisons, thus viable solely for tasks with limited policy diversity—this work embraces the maximum entropy (MaxEnt) IRL to yield policies of maximal diversity, thereby conferring robustness against the noise and uncertainty endemic to human demonstrations, while simultaneously sustaining adaptability to dynamic driving scenarios (Fig. 1).
The main contributions of the paper are listed as below:
  • Our method leverages the Mamba architecture to efficiently model long-range state-space dependencies in vehicle trajectories, augmented by GNN module that encodes inter-vehicle interactions through spatiotemporal message passing.
  • We integrate a reward function module that learns human-preferential driving policies by maximizing the likelihood of demonstrated trajectories, endowing the model with generalization capabilities in analogous scenarios.
  • The proposed algorithm outperforms state-of-the-art methods on benchmark datasets while demonstrating superior cross-scenario generalization in specialized tests, validating its practical viability for real-world autonomous driving applications.
Figure 1: The illustration of Maximum Entropy Inverse Reinforcement Learning based driving behavior modeling on Cross-Scenario Adaptability.
The rest of paper is organized as follows: Section. II presents the proposed framework, detailing its core modules. Section. III describes the experimental setup, presents comprehensive validation results, and provides in-depth analysis of the findings. Finally, Section. IV concludes the paper with a summary of key contributions and discusses promising directions for future research.
II Methodology
This section presents our proposed framework (Fig. 2), Environment-Aware Mamba Trajectory Predictor with MaxEnt IRL, which integrates sequence modeling with IRL to capture human driver behavior. We first formalize the trajectory prediction problem, then detail our encoder-decoder architecture leveraging Mamba blocks for efficient long-sequence modeling and GATs to encode spatial interactions. Finally, we introduce the MaxEnt IRL module to infer driver reward functions from heterogeneous human demonstrations, enabling more human-like and scenario-adaptive trajectory generation.
Figure 2: The overall structure of the proposed trajectory prediction framework.
II-A Problem Statement
Trajectory prediction is achieved by modeling and analyzing input historical trajectories, utilizing the model to forecast future trajectories. We employ $\tau$ to represent the historical trajectory of the target vehicle, while $\hat{\tau}$ denotes its future trajectory. Note that the analysis in this paper assumes a discrete framework, where a trajectory is defined as a sequence of states, i.e., $\tau=[s_{1},s_{2},\dots,s_{L}]$, with $s_{t}$ representing the state of the vehicle at the $t$-th time step such as position, velocity, yaw angle, acceleration. The discussion of trajectory representation will be extended to include inter-vehicle information expressed through graph-theoretic formalism in Section. II-B. $L$ indicates the length of the given trajectory.
While the use of reinforcement learning for prediction inherently involves uncertain approximations of the environment’s dynamic model, the historical trajectory input, which, even as time series data, implicitly encompasses state-action pairs, i.e., $a_{t}=\Delta x_{t}=x_{t+1}-x_{t}$ where $x_{t}$ is the coordinate position of the target vehicle contained in $s_{t}$. This permits the application of the Markov Decision Process (MDP) for modeling, and allows for the use of IRL to analyze the problem as detailed in Section. II-C.
Our goal is to predict the future trajectory of target vehicle, specifically the coordinate position $\{\hat{x_{t}}\}^{T}_{t=L}$, given historical trajectory $\tau\in\Gamma_{S}$, where $T$ is the final step of the prediction. Furthermore, we aims to implicitly infer a policy $\pi^{*}$ that maximizes the likelihood of generated trajectories $\hat{\tau}\sim\pi^{*}$,$\hat{\tau}\in\Gamma_{T}$, where $\Gamma_{S}$ and $\Gamma_{T}$ are the source and target domain of trajectory data. The potential discrepancy between source and target domains provides the fundamental motivation for investigating our method’s cross-scenario adaptability.
II-B Graph-Based Mamba Predictor
Our prediction module adopts the widely utilized encoder-decoder architecture, wherein the embedded historical trajectory $\tau$ is processed by the encoder to yield a latent representation $h$ that encapsulates the spatiotemporal context, which subsequently serves as the input to the decoder for generating the trajectory $\hat{\tau}$. In the proposed architecture, the Mamba model is exclusively employed as the decoder, a deliberate design choice predicated on the observation that its formidable temporal representational capacity—while advantageous for sequential generation—would render an end-to-end fully Mamba-based encoder-decoder architecture prohibitively challenging to optimize.
The State Space Models map an input signal $x(t)\in\mathbb{R}^{D}$ to the output signal $y(t)\in\mathbb{R}^{D}$ via the latent state $h(t)\in\mathbb{R}^{N}$. For application to a discrete autoregressive input sequence of target vehicle states $(s_{0},s_{1},\dots)$ from each preceding timestep, our Mamba module iteratively updates its hidden state $h_{t}$ at each step $t$ (where $t$ ranges from $L+1$ to $T$):
$h_{t}=\bar{A}h_{t-1}+\bar{B}\hat{s}_{t-1},$ (1)
$\hat{s}_{t}=Ch_{t},$ (2)
where $\bar{A},\bar{B}\text{ and }C$ are discrete trainable matrices. By converting the continuous SSM to the discrete one using the zero-order hold discretization rule, the model becomes a sequence-to-sequence mapping framework. Through Eq. (1) and (2), the predicted sequence of states $\{\hat{s_{t}}\}^{T}_{t=L}$ is recursively computed from each hidden state $h_{t}$, and thus obtaining the future coordinate position $\{\hat{x_{t}}\}^{T}_{t=L}$.
To complete the description of our encoder-decoder architecture, the encoder employs a GRU network while following the methodology of [17, 11] to integrate inter-vehicle information into the latent representation $h_{L}$ through a Graph Attention Network (GAT)[18]. Inter-vehicle relationships are modeled as graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the node set, $\mathcal{E}$ denotes the edge set. Each node represents a traffic participant. For feature representation, let $h^{\nu}\in\mathbb{R}^{d_{f}}$ be the features of node $\nu\in\mathcal{V}$, $e^{\nu,\iota}\in\mathbb{R}^{d_{e}}$ be the features of edge $(\nu,\iota)\in\mathcal{E}$. This design choice is particularly crucial as the relational information between target and surrounding vehicles is typically provided as the node and edge features in the heterogeneous graph.
Within this framework, we implement an attention mechanism that computes aggregation weights across each node $\nu$’s inclusive neighborhood $\tilde{\mathcal{N}}(\nu)=\mathcal{N}(\nu)\cup\{\nu\}$, enabling the model to allocate varying importance to different neighbors in the graph. Specifically, we utilize an enhanced attention computation where the attention weights are determined by:
$\alpha_{\nu,\iota}=\frac{\exp\left(\mathbf{a}_{\alpha}^{\top}\gamma\left(\mathbf{W}_{\alpha}[h^{\nu}\|h^{\iota}\|e_{\nu,\iota}]\right)\right)}{\sum_{u\in\tilde{\mathcal{N}}(\nu)}\exp\left(\mathbf{a}_{\alpha}^{\top}\gamma\left(\mathbf{W}_{\alpha}[h^{\nu}\|h^{u}\|e_{\nu,u}]\right)\right)},$ (3)
where $\mathbf{a}_{\alpha}$ and $\mathbf{W}_{\alpha}$ are learnable parameters of GATs; $\gamma(\cdot)$ denotes the Leaky ReLU activation function (with negative slope $0.2$), providing non-linear transformation; $[h^{\nu}\|h^{\iota}][e_{\nu,\iota}]$ indicates concatenation of features. The hidden state of the target vehicle node with graph considered is updated through:
$h^{\prime\nu}=b+\sum_{\iota\in\tilde{\mathcal{N}}(\nu)}\alpha_{\nu,\iota}\mathbf{W}_{\alpha}h^{\iota},$ (4)
where $b$ is a learnable bias term.
II-C Mamba-based MaxEnt IRL
As delineated in Section. II-A, given $N$ observed trajectories $\Gamma=\{\tau_{i}\}^{N}=\{s_{1},s_{2},\dots,s_{L}\}_{N}$, which can be transformed into a set of human demonstrations $\mathcal{D}=\{(s_{t},a_{t})\}^{L\times N}_{t=1}$, our principal objective revolves around the estimation of the underlying reward function, the accurate reconstruction of which would enable the emulation and prediction of nuanced human driving behaviors with enhanced fidelity[19].
Assume that the reward function of the driver is denoted by $R$ parametrized by $\theta_{\text{RF}}$. We employ a neural network with activation layers to fit the reward function for better performance instead of a linear structured reward function in [14]. According to the principle of MaxEnt IRL [12], we learn the reward function by minimizing the negative log-likelihood of the human trajectories, as shown in Eq. (5).
$\mathcal{L}_{\text{RF}}=-\mathbb{E}_{\tau\sim\pi^{*}}\left[\log P(\tau|\theta_{\text{RF}})\right],$ (5)
where $\pi^{*}$ is the human strategy within the human demonstration. The trajectory likelihood $P(\tau)$, governed by the maximum entropy principle, is proportional to the exponential of the trajectory’s cumulative reward. To elaborate with precision, the likelihood of a trajectory $\tau$ necessitates normalization through division by the partition function $Z(\theta_{\text{RF}})$—a quantity explicitly defined in Eq. (6). Here with a bit abuse of notation, the summation $\sum_{\tau}$ enumerates over the entire space of feasible human trajectories, while $R(\tau|\theta_{\text{RF}})$ encapsulates the cumulative rewards associated with a specific trajectory $\tau$ parameterized by $\theta_{\text{RF}}$. This normalization not only guarantees the requisite properties of a probability measure but also introduces a logarithmic regularization term during optimization, thereby mitigating the risk of overfitting to the human demonstrations.
$\begin{split}P(\tau|\theta_{\text{RF}})&=\frac{1}{Z(\theta_{\text{RF}})}\exp\left(\sum_{t=1}^{L}R(s_{t},a_{t}|\theta_{\text{RF}})\right)\\ &=\frac{\exp\left(\sum_{t=1}^{L}R(s_{t},a_{t}|\theta_{\text{RF}})\right)}{\sum_{\tau}\exp\left(R(\tau|\theta_{\text{RF}})\right)}.\end{split}$ (6)
In practical applications, computing the partition function $Z$ is generally infeasible since the trajectory space is typically extremely large. To circumvent this computational impasse, one may resort to importance sampling [19] techniques, whereby $Z$ is approximated through Monte Carlo estimation using $M$ samples from $\Gamma$:
$\hat{Z}(\theta_{\text{RF}})\approx\frac{1}{M}\sum_{t=1}^{M}\frac{e^{R(s_{t},a_{t}|\theta_{\text{RF}})}}{\pi^{*}(a_{t}|s_{t})}.$ (7)
When the optimal policy $\pi^{*}$ remains unspecified and the forward derivation of agent policy is beyond the scope of current investigation, we adopt the simplifying assumption of uniform exploration, where $\pi^{*}(a|s)$ is treated as constant proportional to the trajectory length.
To prevent overfitting of the reward function, a regularization term is added to constraint the norm of reward function parameter $\theta_{\text{RF}}$, yielding the final instantiation of the reward function loss:
$\mathcal{L}_{\text{RF}}=-\mathbb{E}_{\tau\sim\pi^{*}}\left[\log P(\tau|\theta_{\text{RF}})\right]+\lambda\|\theta_{\text{RF}}\|^{2}.$ (8)
The Trajectory Prediction Module (TPM) integrates all components described in Section. II-B, namely the graph-based encoder and Mamba decoder. The TPM updates its parameters by simultaneously optimizing the mean squared error of the predicted coordinates, while also maximizing the cumulative reward of the predicted trajectory in accordance Eq. (5):
$\mathcal{L}_{\text{TPM}}=$ $\frac{1}{T-L}\sum_{t=L}^{T}\left\|x_{t}-\hat{x}_{t}\right\|^{2}$ (9)
$-\gamma\mathbb{E}_{{\hat{\tau}}\sim\pi^{*}}\left[\log P({\hat{\tau}(\theta_{\text{TPM}})|\theta_{\text{RF}}})\right],$
where $\hat{\tau}(\theta_{\text{TPM}})$ is used to clarify the reliance of predicted trajectory and the parameter of the TPM. This integrated approach enables the graph-based Mamba predictor to generate trajectory predictions while the reward function module ensures likelihood maximization of these trajectories. The complete workflow is presented in Algorithm. 1.
Algorithm 1 Mamba predictor with Maxent IRL
0:  Human trajectories $\Gamma=\{\tau_{i}\}^{N}$, $\theta_{\text{TPM}}$, $\theta_{\text{RF}}$, Heterogeneous Graph $\mathcal{G}=\left(\{\mathbf{X}^{t}\}_{t\in\mathcal{T}_{v}}\{\mathbf{E}^{r},\text{Edge Index}^{r}\}_{r\in\mathcal{T}_{e}}\right)$
0:  Optimized trajectory prediction module and reward function parameters $\theta_{\text{TPM}}^{*}$, $\theta_{\text{RF}}^{*}$
1:  Initialize $\theta_{\text{TPM}}$, $\theta_{\text{RF}}$
2:  for epoch $\leftarrow$ 1 to $N$ do
3:     Obtain $h_{L}$ from $\tau$ using encoder
4:     Encode environmental information $\mathcal{G}$ into $h_{L}$ via GAT
5:     Decode $h_{L}$ using Mamba to generate:  - The predicted trajectory $\hat{\tau}$  - Future coordinates $\{\hat{x}_{t}\}_{t=L}^{T}$
6:     Compute reward values from trajectores for maximum likelihood in both $\mathcal{L}_{\text{TPM}}$ and $\mathcal{L}_{\text{RF}}$:  - Historical trajectory rewards $\{r_{i}\}_{t=1}^{L}=R(\tau|\theta_{RF})$  - Predicted trajectory rewards $\{r_{i}\}_{t=L}^{T}=R(\hat{\tau}|\theta_{RF})$
7:     Loss calculation by Eq. (8) and (9):
$\mathcal{L}_{\text{TPM}}\leftarrow$ $\frac{1}{T-L}\sum_{t=L}^{T}\left\|x_{t}-\hat{x}_{t}\right\|^{2}$
$-\gamma\mathbb{E}_{{\hat{\tau}}\sim\pi^{*}}\left[\log P({\hat{\tau}(\theta_{\text{TPM}})|\theta_{\text{RF}}})\right],$
$\mathcal{L}_{\text{RF}}\leftarrow-\mathbb{E}_{\tau\sim\pi^{*}}\left[\log P(\tau|\theta_{\text{RF}})\right]+\lambda\|\theta_{\text{RF}}\|^{2}.$
8:     $\theta_{\text{TPM}}^{*}\leftarrow\theta_{\text{TPM}}$
9:     $\theta_{\text{RF}}^{*}\leftarrow\theta_{\text{RF}}$
10:  end for
III Evaluation and Results
The efficacy of the proposed framework has been validated through an extensive suite of experiments, encompassing diverse driving scenarios and performance metrics. To ensure a holistic assessment, we first conduct a systematic comparison against state-of-the-art baselines across benchmark datasets. Subsequently, we present detailed ablation studies and cross-scenario test to dissect the contributions of individual components, thereby quantifying their impact on overall performance.
Datesets. Our experimental evaluation leverages two large-scale trajectory datasets—the rounD [20] and inD [21] datasets—which comprise vehicular movement patterns captured at 25Hz across diverse German roundabout and intersection scenarios(Fig. 3). These datasets, having undergone preprocessing procedures outlined in [10], were subsequently downsampled by a factor of 5 to achieve a temporal resolution of 0.2s. The observation window was deliberately constrained to 3s ($L$ = 15 timesteps), while the prediction horizon was extended to 5s ($N$ = 40 timesteps). Through this meticulous curation process, the final datasets yield 29,248 and 3,738 high-fidelity samples for rounD and inD, each sample representing a complete spatiotemporal graph sequence. To ensure robust model development and evaluation, the data was partitioned in a stratified manner, with 80% allocated for training, 10% for validation, and the remaining 10% reserved for testing purposes.
Figure 3: Environment and trajectory notation of urban intersection scenario at Heckstrasse.
Metrics The algorithm’s predictive performance was evaluated using - distinct metrics (Tab. I), each carefully selected to provide comprehensive assessment criteria. Note that the definition of ADE, FDE, MR, APDE and CR follows [10].
TABLE I: Metrics Details
Metrics Description
ADE
Average Displacement Error: The average Euclidean
distance in meters between the predicted and
ground truth trajectories.
FDE
Final Displacement Error: The Euclidean distance
in meters between the predicted and ground truth
trajectories at the final time step.
APDE
Average Path Displacement Error: The average minimum
L2-norm between thepredicted positions and
ground truth.
MR
Miss Rate: The percentage of predictions where
the predicted trajectory at the final time step
is not within 2.0 m from the ground truth.
CR
Crash Rate: the proportion of predicted trajectories
that result in collisions with obstacles or other
agents under simulated scenarios.
TABLE II: ROUND DATASET VEHICULAR TRAJECTORY PREDICTION METRICS
metric ADE FDE MR APDE
rounD
CA 4.83 16.2 0.95 3.9
CV 6.49 17.1 0.94 4.34
Seq2Seq 1.46 3.66 0.59 0.82
S-LSTM [22] 1.2 3.47 0.56 0.74
CS-LSTM [6] 1.19 3.57 0.6 0.69
GNN-RNN [17] 1.29 3.5 0.59 0.77
mmTransformer [23] 1.29 3.5 0.59 0.77
Trajectron++ [24] 1.09 3.53 0.54 0.59
MTP-GO [10] 0.96 2.95 0.46 0.59
Mixed Mamba [11] 0.93 3.03 0.27 0.57
Ours 0.91 2.78 0.43 0.57
TABLE III: VEHICULAR TRAJECTORY PREDICTION METRICS ON CROSS-SCENARIO ADAPTABILITY
metric ADE FDE MR APDE CR
inD(seen)
GNN-RNN[17] 0.99 2.84 0.39 0.55 0.0023
mmTransformer[23] 0.95 2.48 0.38 0.59 0.0024
AppL[16] 0.78 2.08 0.32 0.48 0.0023
Ours 0.70 1.95 0.30 0.42 0.0023
rounD(Unseen)
GNN-RNN [17] 5.89 13.98 0.95 3.01 0.0038
mmTransformer[23] 5.36 11.79 0.92 2.67 0.0020
AppL[16] 5.46 12.19 0.95 3.97 0.12
Ours 4.72 10.27 0.94 2.42 0.00069
Figure 4: Cross-Scenario Adaptability (CSA) scores by metric visualized in radar chart.
TABLE IV: ABLATION EXPERIMENT RESULTS
Dataset Index Mamba MaxentIRL GNN ADE FDE MR APDE CR
inD $\text{H}_{1}$ 0.71 1.96 0.31 0.43 0.0023
$\text{H}_{2}$ 0.78 2.08 0.32 0.47 0.0031
$\text{H}_{3}$ 0.72 1.96 0.30 0.44 0.0023
$\text{H}_{4}$ 0.83 2.13 0.44 0.49 0.0035
rounD $\text{H}_{1}$ 0.91 2.78 0.43 0.57 0.00067
$\text{H}_{2}$ 1.05 2.98 0.50 0.66 0.0009
$\text{H}_{3}$ 1.01 2.84 0.45 0.63 0.00074
$\text{H}_{4}$ 1.06 3.01 0.51 0.66 0.00087
$\textbf{rounD}^{*}$ $\text{H}_{1}$ 4.72 10.27 0.94 2.42 0.00069
$\text{H}_{2}$ 4.98 10.65 0.94 3.67 0.00069
$\text{H}_{3}$ 8.64 17.94 0.99 5.37 0.0007
$\text{H}_{4}$ 7.03 14.74 0.99 5.01 0.00065
III-A Results Analysis on rounD
The quantitative results are presented in Tab. II. In contrast to Mixed Mamba [11]—a contemporaneous approach that similarly incorporates the Mamba module—the cited method purportedly achieves superior accuracy on the highD dataset compared to the rounD dataset. Our method achieves state-of-the-art performance across key metrics—including average displacement error , final displacement error and average path displacement error, e.g., FDE of 2.78, representing an 8% reduction compared to the Mixed Mamba, thereby demonstrating its robustness in handling high-variability driving behaviors. The discrepancy of Mixed Mamba performance can be attributed to the inherent simplicity of highway trajectory prediction, where vehicular motion patterns are predominantly linear. Furthermore, the highD dataset’s extensive scale likely augments the Mamba module’s performance. However, the aforementioned work fails to extend Mamba’s advantages to roundabout trajectory prediction, a domain that presents significantly greater challenges due to the non-linear and dynamic nature of vehicular interactions. The result shows our proposed method addresses these limitations with a maximum cumulative reward objective. This integration enables our framework to excel in complex urban roundabout scenarios, such as these in Aachen and Alsdorf.
III-B The Effectiveness of Maxent IRL
To rigorously evaluate the generalization capability and Cross-Scenario Adaptability (CSA) of our proposed methodology, we conducted comprehensive experiments involving training on the inD dataset while validating performance across both inD and rounD datasets, with quantitative results tabulated in Tab. III. This experimental paradigm was deliberately designed to probe the model’s transferability between ostensibly similar yet distinct driving environments - while both datasets exhibit comparable microscopic behaviors such as car-following, lane-changing, and turning maneuvers within confined spatial contexts, they diverge in road geometries and traffic dynamics between intersections and roundabouts.
We calculate the CSA scores of different methods across various metrics for quantitative comparative analysis, which is formally defined as:
$CSA=\underbrace{\alpha\cdot M_{\text{known}}}_{\text{Known Scenario}}+\underbrace{M_{\text{unknown}}}_{\text{Unknown Scenario}}-\underbrace{\beta\cdot\mathcal{D},}_{\text{Degradation Penalty}}$ (10)
where:
$M_{\text{known}}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{M_{i}^{\text{known}}-M_{\min}^{\text{known}}}{M_{\max}^{\text{known}}-M_{\min}^{\text{known}}}\right),$
$M_{\text{unknown}}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{M_{i}^{\text{unknown}}-M_{\min}^{\text{unknown}}}{M_{\max}^{\text{unknown}}-M_{\min}^{\text{unknown}}}\right),$
$\mathcal{D}$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{M_{i}^{\text{unknown}}-M_{i}^{\text{known}}}{|M_{i}^{\text{known}}|},$
with $M_{i}$ being the $i$-th evaluation metric (e.g., ADE, FDE). The terms with “$\max$” and “$\min$” subscripts denote the maximum and minimum values of certain scenario respectively across all evaluated methods, respectively, while $\alpha,\beta$ represents the weighting coefficient. Intuitively, a higher CSA value indicates superior cross-scenario adaptability.
The qualitative results are shown in Fig. 4. For interpretability, we set $\alpha=0.8$, $\beta=0.05$ to account for the substantially larger magnitude of performance degradation compared to inter-method variance on individual metrics. Notably, although the CSA values for MR are comparatively lower across all methods—reflecting the inherent challenge of behavioral prediction in unseen scenarios—our approach still demonstrates state-of-the-art performance within this category. More significantly, our method outperforms all baselines in the remaining metrics, achieving approximately 2 times the CSA scores of the second-best method in both APDE and FDE.
By virtue of employing Maxent IRL, which endows the learned policies with substantially enhanced diversity, our method demonstrates superior performance in the inD scenario (Tab. III). Specifically, the proposed approach achieves significantly lower average displacement error and final displacement error compared to the apprenticeship learning method (AppL [16]) that adheres to the maximum margin theory framework. This empirical result substantiates our framework’s enhanced capacity for holistically approximating the underlying distribution of human driving trajectories, demonstrating more faithful reproduction of naturalistic behavioral patterns compared to AppL approaches.
Ablation Study. The ablation experiments for the proposed method were conducted on the inD and rounD datasets, with the results summarized in Tab. IV. Notably, $\textit{rounD}^{*}$ denotes the cross-scenario evaluation where the model is trained exclusively on urban intersection scenarios and is tested on roundabout environments. The extend of degradation of mmTransformer and GNN-RNN methods in transfer scenarios may be attributed to model complexity and GNN-based environmental encoding mechanisms. Generally speaking, more sophisticated temporal modeling architectures, while exhibiting stronger modeling capacity, tend to demonstrate correspondingly poorer generalization capabilities in novel scenarios. Further analysis on the tabulated data reveals that both the MaxEnt IRL module and the GATs module contribute to trajectory prediction accuracy, with the latter exhibiting even more pronounced enhancement. This phenomenon can be attributed to the fact that the graph-based architecture, introduces additional heterogeneous graph data encompassing information that the IRL method doesn’t contain. However, the IRL component does play a pivotal role in bolstering the model’s generalizability across scenarios: in the $\textit{rounD}^{*}$ column in Tab. IV, the absence of the IRL module precipitates an 83% deterioration in ADE and a more than twofold decline in FDE.
IV Conclusion
We proposed the Environment-Aware Mamba Trajectory Predictor with MaxEnt IRL, a novel framework that integrates Mamba-based sequence modeling with IRL to capture human driver behavior. We developed an encoder-decoder architecture combining Mamba blocks for efficient long-sequence dependency modeling and GATs for encoding spatial interactions among traffic agents. The MaxEnt IRL module further enhanced the framework’s adaptability by inferring diverse driver reward functions from heterogeneous demonstrations, enabling human-like and scenario-aware trajectory generation. Experiments on urban intersections and roundabouts demonstrated that our method achieves state-of-the-art prediction accuracy while exhibiting strong generalization to unseen scenarios, owing to its ability to learn transferable reward structures from human data.
Future research directions could explore explicitly deriving agent policies through the learned reward function, which not only has the potential to further enhance the method’s cross-scenario adaptability but also enables more stable training through techniques like importance sampling for approximating the partition function $Z$ with $\pi^{*}$, broadening the practical applicability in real-world autonomous driving systems.